OpenStudy (anonymous):

What is the exact area that lies beneath the curve y=9x^-4

OpenStudy (anonymous):

\[1\le x \le6\]

OpenStudy (anonymous):

\[\int_1^6\frac{9}{x^4}dx\]

OpenStudy (anonymous):

or \[9\int_1^6\frac{dx}{x^4}\] do you know how to find this?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ok use the power rule backwards you have \(x^{-4}\) so the anti derivative will be \(-\frac{x^{-3}}{3}\)

OpenStudy (anonymous):

multiply by that 9 out front and get \(-3x^{-3}=-\frac{3}{x^3}\) plut in 6, plug in one, and subtract

OpenStudy (anonymous):

thanks i got the right answer